HUBBLE SPACE TELESCOPE DAILY REPORT #2316 PERIOD COVERED: 0000Z (UTC) 02/24/99 - 0000Z (UTC) 02/25/99 Daily Status Report as of 056/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/MA1 7772 (The UV-Upturn of Super-Metal-Rich Globular Clusters) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet observations of the super metal rich globular clusters located in galaxy NGC-224 (a.k.a., M-31, Andromeda Galaxy). Clusters were selected with metal enrichment similar to those encountered in the galaxies' nuclear regions, where strong UV fluxes are observed. A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images (i.e., Bias, Internal Flat, K-Spaot, and Dark) were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.3 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.4 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The second iteration of this proposal did not execute due to the safemode event discussed in 2.3. Otherwise, the observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 6699 (The Globular Cluster System of the Dwarf Elliptical System Galaxies NGC-185 and NGC 205) The WF/PC-2 was used to make observations of the globular cluster system around the nucleus of dwarf galaxy NGC-205. The propose is to carry out a general reconnaissance of the properties of the globular clusters. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. The safemode event discussed in 2.3 occurred during the first iteration of this proposal. The second iteration did not execute at all. 1.7 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The safemode event discussed in 2.3 occurred during the first iteration of this proposal. The second iteration did not execute at all. 1.8 Completed WF/PC-2 7497 (WF/PC-2 Observations of Astrophysically Important Visual Binaries) The WF/PC-2 was to be used to make observations of the visual binary variable star system MU-Cas. This system is a famous metal-deficient G dwarf for which accurate masses will lead to the star's helium content with cosmological implications. This proposal did not execute due to the safemode event discussed in 2.3. 1.9 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was to be used to make high galactic latitude, non-scripted, parallel observations. No iteration of this proposal executed due to the safemode event discussed in 2.3. 1.10 Completed WF/PC-2 6791 (The Collisional Debris of NGC 4438) The WF/PC-2 was to be used to make observations of the collisional debris in galaxy NGC-4438. This system is the nearest example of a recent and close high-velocity collision between two large galaxies, and offers a unique opportunity to learn about the physical processes in high-velocity collisions which can transform cluster galaxies. This proposal did not execute due to the safemode event discussed in 2.3. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 0 Successful : 0 2.2 FHST Updates: Scheduled : 5 Successful : 3 Per HSTAR 7016, the full maneuver updates at 056/0335Z and at 056/0338Z failed. 2.3 Operations Notes: Real-time operations continue to be supported by the PRS system. The problem that caused the HST to enter inertial hold safemode (incorrect ephemeris loads) was understood early yesterday, and a plan to return to normal science operations was initiated and successfully implemented. The GEAs were powered on at 055/1302Z, and a new RGA table was uplinked at055/1958Z. This table was for a four gyro configuration with Gyro-3 in the shadow mode. The SIs were all recovered to operational states. New, correct ephemeris tables were successfully uplinked at 055/2050Z. An ARU/PRT re-oriented the HST to an operational attitude. A health and safety load was generated and uplinked at 055/2144Z, and began execution at 055/2146Z. The science intercept SMS loads were received from the STScI and uplinked at 056/0405Z for the DF-224 and at 056/0413Z for the NSSC-1. The NSSC-1 absolute time pointer was set at 056/0415Z. Science was successsfully re-initiated at 056/0500Z. The following proposals detailed above were affected by the safemode entry: 1.4, 1.6, 1.7, 1.8, 1.9, and 1.10. Please see those items for details. Using ROP SR-1, the SSR EDAC error counter was reset once. The SSA transmitters were turned on and off three times per ROP IC-2. SI C&DH errors were reset four times as directed by ROP NS-5. The NSSC-1 status buffer was dumped at 056/0630Z per ROP NS-3. A TTR was generated for loss of signal and no RF downlink due to the safemode event. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH